From GlueXWiki
Jump to: navigation, search

March 10, 2011 FDC meeting


  1. Production
    • Production status (David)
    • Readiness review program (Lubomir, Dave, Bill)
  2. Electronics
    • Status (Fernando, Chris)
    • Foil production (Roger)
    • Other
  3. Engineering (Bill, David)
    • Conductive tape soldering
    • New mounting template
    • Other
  4. Chamber testing
    • "Upside-down" tests (Lubomir)
  5. Other


Participants: Bill, Mark, Dave, Chris, Roger, Simon, Glenn, Eugene, Beni, and Lubomir.


- Dave: Today Mark Stevens and Brian Kross were at Blue Crab explaining how to cut the cathode foil. Two cuts were made using bad foils. Tomorrow they will continue with the foil gluing procedure. All the frames were inspected for mechanical tolerances. Two wire frames (out of 31) were out of specs. In one of them the o-ring groove was too deep and Bill wanted to refurbish it. The test PCB ring is ready to be laminated to the wire frame; the tooling for that is ready, as well. Later we discussed also drilling holes in the clean room for gas, air and other purposes. Brian Krioss was suggesting also venting out the gas from the clean room, but Bill and others say there's no need to do this.

- The readiness review will be on March 31 with Howard Fenker, Eugene Chudakov, Tim Whitlatch, Bruce Lenser, and Brian Kross in the review committee. Bruce wanted to have couple of weeks in advance: the specifications, drawings, and the QA plan. Bill have the assembly drawings ready by the next meeting. Dave has the QA documents. It was not clear what kind of specifications was he interested in; Eugene will talk to Bruce about that. Bill, Dave, and Lubomir discussed tentative agenda: presentations (Overview-Lubomir, Design-Bill, Production-Dave, Recovery procedures-Bill, Tests-Lubomir), Blue Crab visit, and discussions.


- Chris: expects quotation about the parts for the rigid-flexes. Sharing the space in 126: in addition to the FDC test area, Fernando has plans to set-up testing benches, and Bill needs some space too. Lubomir and Fernando will coordinate the use of the space there. Hall A will take their clean tent within a month and they need one tech from us to help. Bill: maybe there are now Hall D techs available to dismount the tent next week; will talk to Tom.

- Roger talked to Allflex about increasing the production rate; they claimed before they can do 1 set/week starting March 25. They will start with 1set/week, but within few months they can do 3-5sets/week, which is what we wanted. Roger also raised the question how to store the foils and Bill will work on this.


- The first samples of soldering with conductive tape were shown at the meeting. The conductivity is good at one side of the connector, but worse at the other with some missing connections. Bill explains it with the non-uniform pressure, since the tool was used without conductive rubber pad. Bill is going to improve the device and the soldering will be tested again.

- Bill accepts requirements for the mounting template. Lubomir: we need lifting frame for the chamber package and Bill proposed to use the same template as a lifting frame. Bill is working also on tools for the ultrasonic cleaner.

Chamber testing

- The gain problem with the "upside-down" chamber was solved: Lubomir connected the HV cables in a wrong way with all wires, sense and field, at positive high HV, i.e. the chamber was operated as proportional chamber. After fixing this, the top/bottom asymmetry still remains and goes with the gravity. Again we discussed different possibilities: pressure difference, sagging, frame deformation. We also discussed the gas flow in the chamber and the fact that there's limited circulation between the cathode and ground planes.